OpenStudy (sedatefrog712):

Medal and or a fan Part A: Divide (10x4y3 + 5x3y2 - 15x2y - 25x2y4) by -5x2y. Show your work, and justify each step. Part B: How would your answer in Part A be affected if the x2 variable in the denominator was just an x? Part C: What is the degree and classification of the polynomial you got in Part A?

OpenStudy (sedatefrog712):

@robtobey

OpenStudy (sedatefrog712):

@zepdrix

zepdrix (zepdrix):

:o

OpenStudy (sedatefrog712):

what

zepdrix (zepdrix):

\[\Large\rm \frac{10x^4y^3+5x^3y^2-15x^2y-25x^2y^4}{-5x^2y}\]So we need to do this division, yah?

OpenStudy (sedatefrog712):

yes sir

zepdrix (zepdrix):

\[\Large\rm =\frac{10x^4y^3}{-5x^2y}+\frac{5x^3y^2}{-5x^2y}-\frac{15x^2y}{-5x^2y}-\frac{25x^2y^4}{-5x^2y}\]We can separate the terms like this. And do the division piece by piece. Should work out ok. :o

zepdrix (zepdrix):

Or we can do factoring in the numerator if you're more comfortable with that.

zepdrix (zepdrix):

Yah let's do that actually.. that makes more sense. Factoring.

zepdrix (zepdrix):

Err no factoring would be simply doing the division anyway... blahhh how you wanna do this? :d

OpenStudy (sedatefrog712):

which is easiest

zepdrix (zepdrix):

\[\Large\rm =\color{orangered}{\frac{10x^4y^3}{-5x^2y}}+\frac{5x^3y^2}{-5x^2y}-\frac{15x^2y}{-5x^2y}-\frac{25x^2y^4}{-5x^2y}\]What do you get for this first term? Divide the 10 and -5. Divide x^4 by x^2. (recall that when you divide terms with similar bases, you `subtract` the expoennts).

OpenStudy (sedatefrog712):

-2x^2y

zepdrix (zepdrix):

10 divide -5 gives us -2. x^4 divide x^2 gives us x^(4-2) which is x^2. y^3 divide y gives us y^(3-1) which is y^2.\[\Large\rm \color{orangered}{-2x^2y^2}\]Mmmm ok very good! :) Do that same process with the other terms.

zepdrix (zepdrix):

\[\Large\rm =\color{orangered}{-2x^2y^2}+\frac{5x^3y^2}{-5x^2y}-\frac{15x^2y}{-5x^2y}-\frac{25x^2y^4}{-5x^2y}\]

OpenStudy (sedatefrog712):

-1x^1y

OpenStudy (sedatefrog712):

-3xy -5xy^3

zepdrix (zepdrix):

In the last two terms keep in mind that we're subtracting these negative values, so it will change them to positive, yes? \[\Large\rm =-2x^2y^2-xy-\frac{15x^2y}{-5x^2y}-\frac{25x^2y^4}{-5x^2y}\]

zepdrix (zepdrix):

\[\Large\rm \frac{x^2}{x^2}\ne x\]

OpenStudy (sedatefrog712):

ok so -3 -5y^3

zepdrix (zepdrix):

\[\Large\rm =-2x^2y^2-xy-(-3)-(-5y^3)\]Mmm ok good. Which simplifies a little bit.

OpenStudy (sedatefrog712):

By combining like terms?

zepdrix (zepdrix):

No. No like-terms here. Just combine the negatives.

OpenStudy (sedatefrog712):

ok so its xy^4

OpenStudy (sedatefrog712):

I that correct @zepdrix

zepdrix (zepdrix):

What..? I don't understand what you did...

OpenStudy (sedatefrog712):

I combined the negatives

zepdrix (zepdrix):

There are no like-terms. You shouldn't be combining anything. Just simplify the negative signs. Example: \(\Large\rm -(-4)=+4\)

OpenStudy (sedatefrog712):

ok so -2x^2y^2 -xy + 3 + 5y^3

zepdrix (zepdrix):

Yes.

OpenStudy (sedatefrog712):

o ok so is it the same thing with part B

zepdrix (zepdrix):

For part B, If we had divided by x instead of x^2, we would be dividing by `one less power of x`. So each term would end up with `one more power of x`.

OpenStudy (sedatefrog712):

o ok I see thank you so much man

OpenStudy (sedatefrog712):

btw I hope you have a good Shabbat my Jewish brother